A sample of final exam score is normally distributed with a mean equal to 20 and a variance equal to 25

(a) what percentage of scores is between 15 and 25

(b) what raw score is the cutoff for the top 10% of scores

(c) what is the probability of a score less then 27

(d) what is proportion below 13
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 20
std deviation   =σ= 5

a)

probability = P(15<X<25) = P(-1<Z<1)= 0.8413-0.1587= 0.6826

b)

for top 10th percentage values are at 90th percentile:

for 90th percentile critical value of z= 1.28
therefore corresponding value=mean+z*std deviation= 26.40

c)

probability = P(X<27) = P(Z<1.4)= 0.9192

d)

probability = P(X<13) = P(Z<-1.4)= 0.0808